OpenStudy (anonymous):

Find the Solution to the following system of equations -2x+5y=8 2x+2y=6

OpenStudy (anonymous):

-2*x+5*y=8 simplifies to -2*x+5*y-8=0 so hard huh?

OpenStudy (anonymous):

you can take and combine like terms so the -2x and +2x will cancel. then just add up your y's so you will have 7y =14 y= 7 then just plug back in 2x+2(7)=6 2x=-8 x=-4 to check yourself, go 2(-4)+2y=6 -8+2y=6 2y=14 y=7

OpenStudy (anonymous):

also, you need to have sweg.